Q # 2: Consider the following information for an unlevered firm U:
EBIT = BDT 1.6 million annually
Unlevered value VU = BDT4 million
Tax rate = 25%
Cost of debt = 12%
A levered firm L in the same business risk class has a debt-to-equity ratio of 0.5. Use the MM propositions to determine:
a. The after-tax cost of equity for firms U and L.
b. The after-tax WACC for both firms.
1
Expert's answer
2020-12-09T19:43:13-0500

a. Let Unlevered cost of capital=12%


"VU=\frac{EBIT\times(1-T)}{ Unlevered cost of capital}=\frac{1.6\times0.06484746\times(1-0.25)}{0.12}=0.65"


"VL=VU+borrowing\times T=0.65+(0.65\times50)\times25=0.65+0.08=0.73"


b. "WACC=\frac{E}{V}\times Re+\frac{D}{V}\times Rd(1-T)"

"WACCu=\frac{E}{V}\times Re+\frac{D}{V}\times Rd(1-T)= \frac{0.65}{0.65}+\times12+0=12"

"WACCv=\frac{E}{V}\times Re+\frac{D}{V}\times Rd(1-T)= \frac{0.65}{0.73}\times 12+\frac{0.325}{0.73}\times 12(1-0.25)="14.69